OpenStudy (babynini):

Solve the given equation tan^2(theta)-2sec(theta)=2

OpenStudy (babynini):

@freckles I love you <3 just so you know xD I've gotten to sec(theta)=-1/2 and sec(theta)=1

OpenStudy (babynini):

But on the unit circle secant never = -1/2 and it equals 1 at 0

OpenStudy (babynini):

@peachpi :)?

OpenStudy (anonymous):

Pi and -pi work, but they're certainly not the only answers. Positive and negative 1.231, 5.052, and 7.514 work as well.

OpenStudy (babynini):

how did you get those?

OpenStudy (babynini):

does pi(n) where n is an interger work?

OpenStudy (anonymous):

tan² x -2 sec x = 2 sec² x - 1 -2 sec x = 2 sec²x - 2 sec x - 3 = 0 (sec x - 3)(sec x + 1) = 0 sec x = 3, sec x = -1

OpenStudy (babynini):

ooh crud. I added two instead of subtracting it

OpenStudy (babynini):

oh wow I messed up all over the place.

OpenStudy (babynini):

what did you do with the 2secx?

OpenStudy (babynini):

oh I see 2secx-sec^2x-2 = 0

OpenStudy (anonymous):

I factored. and you have them reversed in the equation you just wrote

OpenStudy (babynini):

What's the correct way?

OpenStudy (anonymous):

sec²x - 2 sec x - 3 = 0

OpenStudy (babynini):

ah right. so but secant never = 3 on the unit circle..

OpenStudy (babynini):

and = -1 and pi

OpenStudy (anonymous):

yes, pi is a solution. For sec x = 3, the solution is x = arcsec 3

OpenStudy (babynini):

ah k. So for me with my little scientific calculator how do I do arc sec 3

OpenStudy (babynini):

sec = 1/cos but with the inverse I want to be careful o.o

OpenStudy (anonymous):

sec x = 3 cos x = 1/3 x = arccos (1/3)

OpenStudy (babynini):

oh! :)

OpenStudy (anonymous):

that'll give you an angle in the 1st quadrant. Subtract it from 360 to get the angle in the 4th quadrant.

OpenStudy (anonymous):

or rather 2pi

OpenStudy (babynini):

does it work to add two pi to that? or just subtracting it from 2pi?

OpenStudy (babynini):

the answers they give are kpi = we have this one! 0.72+2kpi = we don't have this. 5.56+2kpi = and we dont' have this either.

OpenStudy (anonymous):

adding 2pi just gives the same angle you already found. Subtracting x from 2pi gives the angle in the 4th quad. Where did these answers come from?

OpenStudy (babynini):

aah I see the answer part of the book I'm doing the odd problems in my textbook which gives the answers in the back.

OpenStudy (babynini):

oh crap. Just a moment.

OpenStudy (anonymous):

I'm getting 5.05 and 1.23

OpenStudy (babynini):

Sorry, yeah. I was looking at 7 instead of 5. the answers are (2k+1)pi, 1.23+2kpi, 5.05+2kpi

OpenStudy (anonymous):

ok. you know how to get those?

OpenStudy (babynini):

um 1.23, 5.05 came from theta = arcos(1/3) and pi came from sec(theta)=-1

OpenStudy (babynini):

and I understand the adding 2pi(k)

OpenStudy (babynini):

I'm not sure why the pi one is written as (2k+1)pi ?

OpenStudy (anonymous):

That's the same as π+2πk. They factored the pi out

OpenStudy (babynini):

ah ok I see.

OpenStudy (babynini):

haha I told myself I would keep doing these until I get the answer entirely on my own..but going like this it seems that won't be anytime soon =.= K, this one makes sense now though :)

OpenStudy (anonymous):

ok great. Maybe take a break if you don't have a deadline. These definitely take time to get used to

OpenStudy (babynini):

Final on Friday..hah but I may move onto studying another part.

OpenStudy (anonymous):

ah.. I don't envy you. good luck

OpenStudy (babynini):

yeah xD thanks.

OpenStudy (babynini):

Precalc 2 man. phew.

OpenStudy (babynini):

For you it's probably no biggy xP

OpenStudy (anonymous):

yeah give it a few years, you'll either never use it, or wonder how you lived without knowing it. There's no in-between

OpenStudy (babynini):

I fear it may be the latter D: Just because i despise it so much, i'm bound to end up being a math major lol

OpenStudy (babynini):

Welp, moving onto question next. Thanks for all your help :)